A small funnel in the shape of a cone is being emptied of fluid at the rate of 12 cubic

centimeters per second. The height of the funnel is 20 centimeters, and the radius of the top is 4 centimeters. How fast is the fluid level dropping when the level stands 5 centimeters above the vertex of the cone?




1
Expert's answer
2021-05-24T19:17:24-0400

Ans:-

We are given information about the rate of change of volume of water, so we are given that "\dfrac{dV} {dt} = −12" . Note that the formula for volume is given in terms of r and h, but we only want "\dfrac{dh }{dt}" . We need a relationship between r and h... You should see similar triangles (Draw a line right through the center of the cone. This, and the line forming the top radius are the two legs. The outer edge of the cone forms the hypotenuse).

"\dfrac{radius of top }{radius of water level} = \dfrac{overall height} {height of water} ⇒ \dfrac{4} {r} = \dfrac{20} {h}"


so that "r = \dfrac{h }{5}" . Substituting this into the formula for the volume will give the volume in terms of h alone:

"V =
\dfrac{1}
3
πr^2h =
\dfrac{1}
3
π(
\dfrac{h}
5
)^
2h =
\dfrac{1}
{75}
πh^3"


Now,

"\Rightarrow \dfrac{dV }{dt} = \dfrac{3π }{75 }h ^2 \dfrac{dh} {dt}"


and we know "\dfrac{dV} {dt} = −12, h = 5," So


"\Rightarrow \dfrac{dh}{
dt }=
\dfrac{−12}
π"